OpenStudy (anonymous):

what are the tranformations of this function? 2log(x+3)

OpenStudy (anonymous):

To find the inverse function you need, practically, to exchange x and y. So it would be like this: x = -2log base 5 (y + 3) + 1 And then, isolate y: log base 5 (y + 3) = (1 - x)/2 y + 3 = 5^((1 - x)/2) y = inverse function = 5^((1 - x)/2) - 3 Transformations: multiplying by -2 the graph will have to be put upside down with a vertical stretch (2x). Adding 1, it will go up by 1 unity. Adding 3 to x, it will go left by three unities.

OpenStudy (anonymous):

I am graphing this function. I know that they transformations are 1. vertical strech by 2 and left 3...do I multiply the 2 from the strech with the x's or y's from my coordinate points

OpenStudy (anonymous):

correct x

OpenStudy (anonymous):

so i multiply it by the x's?

OpenStudy (anonymous):

is there a way you can show me what the graph will look right?

OpenStudy (anonymous):

can you show me what the graph will look like